R 100 000 is deposited into an account now that pays interest at a rate of 8% p.a. so that an amount of money can be withdrawn from the account every six months in perpetuity starting one year from now. If it is decided to increase the value of the withdrawals, from the second withdrawal onwards, at a rate of 6% p.a., then the value of the first withdrawal (rounded to the nearest cent) made one year from now is equal to
1
Expert's answer
2020-05-20T19:23:38-0400


this is a perpetual annuity

"100 000\times0.04+104 000\times0.03=4000+3120=7120"